### v1

# Shuffle Mamba: State Space Models with Random Shuffle for Multi-Modal Image Fusion

Abstract
Multi-modal image fusion integrates complementary information from different modalities to produce enhanced and informative images. Although State-Space Models, such as Mamba, are proficient in long-range modeling with linear complexity, most Mamba-based approaches use fixed scanning strategies, which can introduce biased prior information. To mitigate this issue, we propose a novel Bayesian-inspired scanning strategy called Random Shuffle, supplemented by an theoretically-feasible inverse shuffle to maintain information coordination invariance, aiming to eliminate biases associated with fixed sequence scanning. Based on this transformation pair, we customized the Shuffle Mamba Framework, penetrating modality-aware information representation and cross-modality information interaction across spatial and channel axes to ensure robust interaction and an unbiased global receptive field for multi-modal image fusion. Furthermore, we develop a testing methodology based on Monte-Carlo averaging to ensure the model’s output aligns more closely with expected results. Extensive experiments across multiple multi-modal image fusion tasks demonstrate the effectiveness of our proposed method, yielding excellent fusion quality over state-of-the-art alternatives. Code will be available upon acceptance.
Introduction
Multi-modal image fusion, a fundamental task in computer vision, involves extracting and integrating valuable information from images of the same scene captured by different imaging modalities. This task aims to create a single composite image with a more comprehensive and informative representation, with typical applications including pan-sharpening and medical image fusion (MIF). In the context of pan-sharpening, satellites are limited by sensors, which can only capture low-resolution multi-spectral (LRMS) and panchromatic (PAN) images. Specifically, PAN images offer superior spatial details but limited spectral resolution, while MS images provide abundant spectral resolution but lack spatial clarity. Through integrating the complementary information from both MS and PAN images into a composite representation, we can achieve an effective balance between spatial and spectral resolution. Analogously, in the realm of MIF, various imaging technologies capture distinct types of information. For instance, Computed Tomography (CT) images deliver detailed insights into bones and high-density tissues, while Magnetic Resonance Imaging (MRI) offers higher-resolution images with rich soft tissue details. In virtue of this complementary information from various modalities, MIF can overcome the limitations of single-modality images, resulting in a more comprehensive and detailed representation for modern medical diagnosis.
Figure 1: Visualization of Effective Receptive Fields (ERFs) for Conv, Self-Attention and our Method. A larger ERF is indicated by a more extensively distributed dark area.
In recent years, the prosperous advancement of deep neural networks (DNNs) has led to the development of numerous DNN-based multi-modal image fusion methods. In pan-sharpening, the pioneering work PNN (Masi et al. 2016) employed a simple three-layer neural network to achieve remarkable results which was previously deemed unattainable, highlighting the superior learning capabilities of deep learning. After that,increasingly complex and deeper architectures have emerged (Zhou et al. 2022b; Yang et al. 2023), delivering excellent visual performance. Despite these advancements, existing multi-modal fusion methods face common limitations. Convolutional neural network (CNN)-based approaches often struggle to establish global receptive fields. While transformers address this issue through self-attention mechanisms, they introduce significant challenges related to quadratic computational complexity. Nowadays, structured state-space models have gained considerable attention for their computational efficiency and principled ability to model long-range dependencies. However, their selective scanning mechanism can introduce biased priors when processing 2D images. To overcome these challenges, it is reasonable for us to design a novel sequential scanning method with its application framework.
Our Motivation. Global modeling capabilities are essential in image restoration tasks because one of the core aspects of image restoration is to find useful information within the image to compensate for the missing data in the current patch. For a long time, CNN and ViT have been dominant architectures in computer vision, each with distinct advantages and limitations. CNNs are constrained by local receptive fields, which hinder their ability to model long-range dependencies. In contrast, ViT uses the self-attention mechanism to access a global receptive field but are burdened by quadratic computational complexity. Recently, structured state-space models have demonstrated enhanced capabilities in capturing long-term dependencies in sequence data while maintaining linear time complexity. Notably, Mamba has achieved significant strides in reducing inference latency and improving overall performance through selective state spaces and hardware-aware algorithms. With the introduction of vmamba (Liu et al. 2024) and Vision Mamba (Zhu et al. 2024), there has been a growing interest in applying state-space models to visual tasks.
However, most current Mamba-based methods employ unidirectional SSMs, which endure certain limitations in their receptive field range due to their scanning approaches. Specifically, the receptive field is large for the initial portion of input patches but becomes significantly smaller for the latter portion, potentially compromising the model’s global modeling capabilities. Moreover, dissimilar to the dependencies between token orders in language modeling, the non-causal nature of 2-D spatial information in images presents a considerable challenge for simple sequential scanning methods. Traditional strategies, such as flattening image patches and scanning them sequentially, may introduce biased local 2D dependencies, thereby undermining the model’s ability to effectively identify spatial relationships.
To address these challenges, we propose a new sequence scanning method called Random Shuffle Scanning. Figure 1 shows the visualization of the effective receptive field (ERF) of three methods. Compared to the Conv method, the Self-Attention method’s advantage lies in achieving a global receptive field. However, our method provides an overall unbiased global receptive field, enabling the network to discard fixed local priors and focus more effectively on what needs to be learned. For the sequential input image patches, we first apply position encoding, followed by a random shuffle of the patches before they are processed by the Mamba block for long-range dependency modeling. The random shuffle approach eliminates the preconception between local and global 2D dependencies in mathematical expectation, enabling the model to access an unbiased prior and ultimately establish a more consistent and global receptive field. Recognizing that shuffling the spatial position of the image patches may disrupt semantic information, we implement a corresponding inverse transformation to accurately restore the sequence order of the input patches after passing through the Mamba block. This information-preserving transformation pair underpins the key components of our model: Random Mamba Block, Random Channel Interactive Mamba Block, and Random Modal Interactive Mamba Block. Inspired by Dropout, we use Monte Carlo averaging to approximate the expected output, ensuring that the actual output at test time closely matches the anticipated results.
Our contributions can be summarized as follows: 1) We design the Shuffle Mamba framework, where the random shuffle operation in the key component provides an expected unbiased global receptive field without increasing any parameters. 2) We develop a specific strategy for training and testing this framework. During training, each input is independently scanned using a random shuffle operation. During testing, we use Monte Carlo averaging to estimate the output of each Mamba block. 3) Extensive experiments on two prominent multi-modal image fusion tasks demonstrate that our method accomplishes excellent performance in both quantitative assessments and visual quality.
Related Works
State Space Model
The State Space Model (SSM), originating from control theory, has found extensive applications in deep learning due to its remarkable ability to model long-range dependencies. Initially, the S4 (Gu, Goel, and Ré 2021) model introduced the concept of the SSM, effectively reducing the computational and memory requirements associated with state representation while enabling global information modeling. Building upon S4, the S5 (Smith, Warrington, and Linderman 2022) model features MIMO structure and efficient parallel scanning strategy to enhance performance without significantly increasing computational demands. Furthermore, the H3 (Mehta et al. 2022) model further refines these methods, achieving competitive performance and efficiency comparable to the transformer in language modeling tasks.
Figure 2: The architecture of the proposed Shuffle Mamba framework.
Recently, Mamba (Gu and Dao 2023) has significantly improved inference speed and performance metrics through selective state spaces and hardware-aware algorithms. The introduction of Vmamba (Liu et al. 2024) and Vision Mamba (Zhu et al. 2024) has brought attention to the application of SSM in vision tasks. However, most existing SSM-based vision models (Zhu et al. 2024; Ma, Li, and Wang 2024; Zheng and Wu 2024) employ a fixed scanning strategy, which may introduce preconceptions, particularly in low-level vision tasks (Xiao et al. 2023). Specifically, this fixed order of choosing image patches can cause the model to gradually dismiss previous input sequences while processing the current patch, thereby compromising its capability to model global information. To alleviate this challenge, Vmamba (Liu et al. 2024) introduced the CSM, which scans image pixels from various directions such as top-left, bottom-right, top-right, and bottom-left. Building on this idea, RSM (Zhao et al. 2024) devised the OSSM to flatten image patches into sequences in eight directions, enhancing the network’s capability to capture and model large-scale spatial features. Additionally, LocalMamba (Huang et al. 2024) applies the windowed selective scan to ensure a harmonious integration of global and local visual cues. RSMamba (Chen et al. 2024) incorporates dynamic multi-path activation mechanisms to model non-causal data. However, these methods do not simultaneously consider the integrity of the image’s structure and the randomness of the pixels during scanning processing, and are essentially still special fixed strategies. To address this limitation, we introduced the Shuffle Mamba framework, which includes random shuffle and inverse operations to obtain a global receptive field without bias in expectations.
Method
Preliminaries
Inspired by continuous linear time-invariant (LTI) systems, SSMs exploit an implicit latent state $h(t)\in{\mathbf{R}^{N}}$to map a 1-D sequence $x(t)\in\mathbf{R}$ to $y(t)\in\mathbf{R}$. Specifically, SSMs can be mathematically expressed as an ordinary differential equation (ODE):
${h}^{\prime}(t)=\mathbf{A}h(t)+\mathbf{B}x(t),$ (1)
$y(t)=\mathbf{C}h(t).$ (2)
Where $\mathbf{A}\in{\mathbf{R}^{N\times N}}$ is the evolution matrix, while $\mathbf{B}\in{\mathbf{R}^{N\times 1}}$ and $\mathbf{C}\in{\mathbf{R}^{1\times N}}$ serve as the projection parameters. However, solving these differential equations in a deep learning context can be challenging. The S4 and Mamba models propose introducing a timescale parameter $\mathbf{\Delta}$ to convert continuous parameters $\mathbf{A}$, $\mathbf{B}$ into their discrete counterparts $\bar{\mathbf{A}}$, $\bar{\mathbf{B}}$:
${{h}_{\text{t}}}=\bar{\mathbf{A}}{{h}_{t-1}}+\bar{\mathbf{B}}{{x}_{t}},$ (3)
${{y}_{t}}=\mathbf{C}{{h}_{t}},$ (4)
$\bar{\mathbf{A}}=\mathbf{\exp}\mathbf{(\Delta A)},$ (5)
$\bar{\mathbf{B}}={{\mathbf{(\Delta A)}}^{-1}}(\exp\mathbf{(\Delta A)-I)}\cdot\mathbf{\Delta B}.$ (6)
Finally, the output of the system can be attained through global convolution:
$\bar{\mathbf{K}}=(\mathbf{C\bar{B}},\mathbf{C\bar{A}\bar{B}},...,\mathbf{C{{\bar{A}}^{L-1}}\bar{B}})$ (7)
where $\mathbf{L}$ represents the length of the sequence $x$, $\bar{\mathbf{K}}\in\mathbf{{R}^{L}}$ is a structured convolution kernel.
Network Architecture
The proposed Shuffle Mamba framework consists of three functional components: the Random Mamba block (RM block), the Random Channel Interactive Mamba block (RCIM block), and the Random Modal Interactive Mamba block (RMIM block). The overall workflow is illustrated in Figure 2. Assuming that the input images with different modalities are ${\mathbf{M}_{up}}$ and ${\mathbf{P}_{in}}$, where ${\mathbf{M}_{up}}$ serve as the upsampled ${\mathbf{M}_{in}}$, we first use convolutional layers to project the images into the feature space. Given the limited receptive field of the convolutional layers, which makes capturing global features challenging, we perform patch embedding and send the resulting patches to the RM block for global feature extraction. This process yields the global modality-specific features $\mathbf{F}_{n}^{m}$ and $\mathbf{F}_{n}^{p}$, where $PE=PatchEmbed$:
$\mathbf{F}_{0}^{m},\mathbf{F}_{0}^{p}=PE(\phi({\mathbf{M}_{up}})),PE(\phi({\mathbf{P}_{in}}))$ (8)
$\mathbf{F}_{n}^{m}={{\psi}_{1n}}\cdot\cdot\cdot({{\psi}_{11}}({{\psi}_{10}}(\mathbf{F}_{0}^{m}))$ (9)
$\mathbf{F}_{n}^{p}={{\psi}_{2n}}\cdot\cdot\cdot({{\psi}_{21}}({{\psi}_{20}}(\mathbf{F}_{0}^{p}))$ (10)
The global modality-specific features are then sent to the RCIM block for simple channel information exchange, without introducing additional parameters. The exchanged features continue to be processed by their respective RM block to obtain $\mathbf{F}_{k}^{m}$ and $\mathbf{F}_{k}^{p}$. Next, we use the RMIM block to attain the $\mathbf{F}_{k+j}^{m}$ through the deep fusion of modality features $\mathbf{F}_{k}^{m}$ and $\mathbf{F}_{k}^{p}$. Thus, the final fused image ${\mathbf{H}_{out}}$ can be accessed by reshaping and residual connection:
$\mathbf{F}_{k+j}^{m}={{\theta}_{j}}\cdot\cdot\cdot({{\theta}_{1}}({{\theta}_{0}}(\mathbf{F}_{k}^{m},\mathbf{F}_{k}^{p}),\mathbf{F}_{k}^{p}),\mathbf{F}_{k}^{p})$ (11)
${\mathbf{H}_{out}}=\phi(reshape(\mathbf{F}_{k+j}^{m}))+{\mathbf{M}_{up}}$ (12)
Figure 3: The Random Shuffle Scanning for training.
Figure 4: The Monte-Carlo averaging for testing.
Key Components
Random Shuffle Scanning.
Mamba was originally designed to adapt to the modeling of language sequences. We devised a Random Shuffle Scanning approach to access unbiased local and global dependencies in 2D image while ensuring a global receptive field similar to self-attention. As shown in Figure 3, for sequentially input 2D image patches, we first apply depth-wise convolution for position modeling. The image patches are then randomly shuffled and sent to the Mamba block for long-range dependency modeling. This strategy allows the Mamba block to simulate interactions between adjacent patches with equal probability, enabling the network to learn and model from an unbiased prior effectively. Additionally, since the relative positions of the patches are crucial for reconstructing semantic information, the output image must be accurately aligned with the input based on the inverse shuffle. For this reason, the random shuffle and its inverse operation constitute an information-lossless transformation pair.
Random Mamba Block.
Based on this shuffle-inverse pair, we designed the Random Mamba Block. First, layer normalization is performed on the input feature ${\mathbf{F}_{l-1}}$ to obtain ${{\mathbf{F}^{\prime}}_{l-1}}$ , which is then projected into $\mathbf{x}$ and $\mathbf{z}$ using random shuffle and multi-layer perceptrons (MLPs). In the first branch, $\mathbf{x}$ passes through 1-D convolution layers with SiLU activation to produce $\mathbf{x}^{\prime}$. The SSM is then used to calculate the output $\mathbf{y}$. In the other branch, $\mathbf{z}$ is sent to the activation function to generate the gating for $\mathbf{y}$. Finally, we apply the inverse shuffle and residual connection to get the final output sequence ${\mathbf{F}_{l}}$.
Random Channel Interactive Mamba Block.
In the RCIM block, we refer to the approach from (He et al. 2024a) to achieve lightweight feature interaction between different modalities. We use the split operation to divide modality features $\mathbf{F}_{k}^{m}$ and $\mathbf{F}_{k}^{p}$ into two halves based on the channel dimension, followed by complementary splicing. The exchanged features are then sent to their respective RM Blocks for processing. By repeating these steps, the global modality-specific features are initially fused.
Random Modal Interactive Mamba block.
Motivated by cross-attention, we designed the RMIM block for processing multi-modal image information. In this approach, we project the shuffled sequence features into a shared space and use a gating mechanism to learn complementary information under an unbiased prior, thereby reducing the interference of redundant features on the fusion results. We employ a process similar to RM Block to generate ${\mathbf{y}_{m}}$ and ${\mathbf{y}_{p}}$, and use the input $\mathbf{F}_{l-1}^{m}$ to generate the gating parameter $\mathbf{z}$ for dynamic adjustment of ${\mathbf{y}_{m}}$ and ${\mathbf{y}_{p}}$. The two outputs are then combined and projected, followed by inverse shuffle and reshape operations to align with the input sequence. Finally, the output $\mathbf{F}_{l}^{m}$ of the module can be obtained through depth-wise convolution and feature flattening.
Testing with Monte Carlo averaging
Table 1: Quantitative comparison of pan-sharpening task on three datasets. Bold and underline show the best and second-best values, respectively. $\uparrow$ indicates that the larger the value, the better the performance, and $\downarrow$ indicates that the smaller the value, the better the performance.
WorldView-II Gaofen-2 WorldView-III
[…]
INNformer AAAI’22 41.6903 0.9704 0.0227 0.9514 47.3528 0.9893 0.0102 0.5479 30.5365 0.9225 0.0747 3.1142
SFINet ECCV’22 41.7244 0.9725 0.0220 0.9506 47.4712 0.9901 0.0102 0.5479 30.5901 0.9236 0.0741 3.0798
MSDDN TGRS’23 41.8435 0.9711 0.0222 0.9478 47.4101 0.9895 0.0101 0.5414 30.8645 0.9258 0.0757 2.9581
PanFlowNet ICCV’23 41.8548 0.9712 0.0224 0.9335 47.2533 0.9884 0.0103 0.5512 30.4873 0.9221 0.0751 2.9531
FAME AAAI’24 42.0262 0.9723 0.0215 0.9172 47.6721 0.9898 0.0098 0.5542 30.9903 0.9287 0.0697 2.9531
DISPNet AAAI’24 41.8768 0.9702 0.0221 0.9157 47.4529 0.9898 0.0111 0.5532 30.0426 0.9153 0.0776 3.2620
Pan-mamba arxiv’24 42.2354 0.9729 0.0212 0.8975 47.6453 0.9894 0.0103 0.5286 31.1740 0.9302 0.0698 2.8910
Ours - 42.3428 0.9734 0.0208 0.8840 47.9180 0.9903 0.0097 0.5109 31.4005 0.9327 0.0676 2.8098
Table 2: Evaluation of our method on real-world full-resolution scenes from the WV2 dataset. Bold and underline show the best and second-best values, respectively.
Metric SFIM Brovey GFPCA INNformer SFINet PanFlowNet FAME DISPNet Pan-mamba Ours
$D_{\lambda}$downarrow 0.1403 0.1026 0.1139 0.0995 0.1034 0.0966 0.0951 0.0944 0.0966 0.0941
$D_{S}$downarrow 0.1320 0.1409 0.1535 0.1305 0.1305 0.1274 0.1263 0.1264 0.1272 0.1266
QNR $\uparrow$ 0.7826 0.7728 0.7532 0.7858 0.7827 0.7910 0.7933 0.7938 0.7911 0.7939
Figure 5: Comparative visual experiments of several methods on WV3 datasets
We incorporate stochastic factors into the random shuffle operation, necessitating the marginalization of these factors in the generation of the final fusion. However, the random shuffle method presents a theoretical challenge due to the exponentially large number of potential models, making precise averaging of their predictions infeasible. Drawing inspiration from dropouts (Srivastava et al. 2014), we approximate the expected value of the entire model through layered expectations. Therefore, the computation of the random shuffle during testing can be expressed as follows, where $\mathbf{RM}=RandomMamba$:
$\mathbf{R}{\mathbf{M}^{test}}(x)=\underset{\mathbf{\mathcal{S}}}{\mathop{\mathbb{missing}}{E}}\,[\mathbf{RM}(x,\mathbf{\mathcal{S}})]$ (13)
In fact, estimating $\mathbf{R}{{\mathbf{M}}^{test}}$ based on the aforementioned equation requires enumerating all possible shuffle results, which imposes a significant computational burden. Therefore, we employ Monte Carlo averaging to estimate its expectation:
$\mathbf{R}{\mathbf{M}^{test}}(x)\approx\frac{1}{M}\sum\limits_{i=1}^{M}{\mathbf{RM}(x,{{\mathbf{\mathcal{S}}}_{i}})}$ (14)
Specifically, the input image is independently shuffled ${M}$ times, and then the ${M}$ outputs of $\mathbf{RM}$ are calculated. The mean of these outputs is computed to obtain the final estimate. When ${M}\to\infty$ is present, the Monte Carlo estimator closely approximates the true mean. Figure 4 illustrates the testing process we designed, which significantly reduces the actual testing time by incorporating multiple identical inputs in a mini-batch and utilizing GPUs for parallel computation.
Table 3: Quantitative comparison of MIF task on three datasets. bold and underline show the best and second-best values.
PET CT SPECT
[…]
PSLPT 0.888 0.548 0.373 0.815 0.675 0.502 0.432 0.810 0.850 0.359 0.325 0.933
EMFusion 0.943 0.685 0.783 1.221 1.190 0.552 0.475 1.266 0.885 0.665 0.692 1.212
MSRPAN 1.017 0.581 0.799 1.182 1.319 0.436 0.455 1.261 0.960 0.525 0.560 1.153
SwinFusion 1.642 0.703 0.683 0.725 1.537 0.522 0.545 0.579 1.678 0.744 0.720 0.684
Zero 0.950 0.635 0.774 1.162 0.768 0.320 0.582 1.199 1.046 0.582 0.681 1.180
U2Fusion 0.947 0.460 0.292 0.494 0.309 0.074 0.489 0.042 0.865 0.419 0.696 0.479
CDDFuse 1.481 0.650 0.765 1.227 1.589 0.526 0.530 1.224 0.995 0.786 0.719 1.169
Ours 1.491 0.797 0.741 1.256 1.580 0.546 0.592 1.330 1.470 0.820 0.747 1.240
Figure 6: Comparative visual experiments of several methods on MRI-PET datasets
Table 4: Ablation for our methods on three datasets. The PSNR/SSIM/SAM/ERGAS values on benchmarks are reported. The best results are shown in bold.
Ablation Variant WorldView-II Gaofen-2 WorldView-III
Baseline - 42.3428/0.9734/0.0208/0.8840 47.9180/0.9903/0.0097/0.5109 31.4005/ 0.9327/0.0676/2.8098
Core Operation RM w/o Random Shuffle 42.1460/0.9723/0.0211/0.9042 47.4284/0.9890/0.0102/0.5355 31.3915/0.9329/0.0676/2.8113
RCIM w/o Random Shuffle 42.2443/0.9729/0.0210/0.8919 47.6847/0.9899/0.0099/0.5292 31.3476/0.9325/0.0680/2.8230
RMIM w/o Random Shuffle 42.2136/0.9727/0.0211/0.8978 47.5787/0.9895/0.0100/0.5324 31.0971/0.9295/0.0704/2.9097
Loss Function
In accordance with established norms within this area, the loss function employed in our model for pan-sharpening is the L1 loss. In the MIF task, we use two input images and the fused image to calculate a composite loss function that includes L1 loss, SSIM loss, and gradient loss.
Experiments
Datasets and Benchmark
For the pan-sharpening task, we utilized datasets from WorldView-II (WV2), GaoFen2 (GF2), and WorldView-III (WV3), encompassing a variety of urban and natural scenes. We generated training samples following the Wald (Wald, Ranchin, and Mangolini 1997) protocol in the inaccessible ground truth. We conducted a thorough comparison of our proposed method against classical approaches and deep learning-based methods, including INNformer (Zhou et al. 2022a), SFINet (Zhou et al. 2022b), MSDDN (He et al. 2023), PanFlowNet (Yang et al. 2023), FAME (He et al. 2024b), DISPNet (Wang et al. 2024) and Pan-mamba (He et al. 2024a).
In the MIF task, we employed medical images from the Harvard Medical website, comprising pairs such as MRI-CT, MRI-PET, and MRI-SPECT. For this task, we compared our method with various deep learning-based multi-modal fusion techniques, including PSLPT (Wang, Deng, and Vivone 2024), EMFusion (Xu and Ma 2021), MSRPAN (Fu et al. 2021), SwinFusion (Ma et al. 2022), Zero (Lahoud and Süsstrunk 2019), U2Fusion (Xu et al. 2020), and CDDFuse (Zhao et al. 2023).
Implementation Details
All experiments were conducted using the PyTorch framework on two NVIDIA RTX 3060 GPUs. We trained for 500 epochs with a batch size of 4 for the pan-sharpening task and 200 epochs with a batch size of 1 for the medical image fusion (MIF) task. The network parameters were optimized using the Adam optimizer. The initial learning rate was set to $45e-$ 4, which was subsequently reduced to $85e-$ 8 using the CosineAnnealingLR scheduler over the specified epochs. For the pan-sharpening task, we randomly cropped training set images to obtain LRMS patches of size 32x32 and PAN images of size 128x128. For the MIF task, the training set images were cropped to 256x256.
Comparison with SOTA Methods
Pan-sharpening.
The experimental results on three datasets are presented in Table 1. Reference metrics, including PSNR, SSIM, SAM, and ERGAS (Alparone et al. 2007), are used to evaluate the fusion effect. The results demonstrate that the proposed method outperforms the SOTA methods across all metrics. Notably, in PSNR metric, our method achieves improvements of 0.1047, 0.2727, and 0.2301 dB over Pan-mamba, which has the second-best performance. Better PSNR and SSIM evidence that the fusion results transfers more information from the original image and experiences less distortion. In the qualitative comparison, Figure 5 shows a representative sample from the WV3 dataset. Our method performs better on the MSE graph, indicating that the fusion result is closer to the ground truth. Compared to other methods, our approach achieves a more accurate restoration of spectral and spatial details, highlighting the advantages of our fusion technique.
To further verify the generalization ability of our method in full-resolution scenes, we use three non-reference metrics, including ${{D}_{s}}$, ${{D}_{\lambda}}$ and QNR, to evaluate our method on the full WorldView-II dataset. Table 2 presents the experimental results. Our method outperforms the comparison methods across all results, demonstrating the strong adaptability of Shuffle Mamba in image fusion.
Medical Image Fusion.
The quantitative comparison results of four metrics on the MIF dataset are shown in Table 3. The proposed method performs well across almost all metrics, demonstrating its effectiveness in medical image fusion. In our experimental results, a higher VIF indicates closer alignment with human perception. Improved SCD, Qabf, and SSIM scores suggest that the fused image maintains higher similarity and experiences less distortion compared to the original images. Figure 6 presents a qualitative comparison of several methods on the MRI-PET dataset. Our method exhibits superior visual quality, a finding that is corroborated by the experimental metrics.
Ablation Experiments
To verify the effectiveness of the Random Shuffle operation, we conducted corresponding ablation experiments. The core operations of these experiments involved removing the shuffle operations from the three main modules. Notably, in the experiment where RM-related components were removed, the RM block in the RCIM module was excluded. The experimental results are presented in Table 4. Removing the shuffle operation resulted in a significant decrease in model performance, indicating that the new scanning strategy designed under the Shuffle Mamba framework effectively enhances the quality of multi-modal image fusion.
Due to Monte Carlo averaging, Shuffle Mamba makes trading memory and computational resources possible for improved performance. We studied the relationship between the number of samples, the PSNR index, and resource consumption. We experimented five times for each sample size, calculating the average and standard deviation of diverse metrics. Figure 7 (a) illustrates the PSNR trend, while Figure 7 (b) and (c) details the corresponding memory usage and the time required to process each image. As the number of samples increases, performance and resource consumption rise, enabling a trade-off between performance and efficiency. Additionally, Monte Carlo averaging enhances the theoretical robustness of the random shuffle operation, effectively improving the mean of PSNR while reducing the variance of the fusion result. More experimental results can be found in the supplementary materials.
Figure 7: The performance and costs in testing time and memory when choosing different numbers of samples.
Conclusion
In this paper, we replace the fixed scanning strategy used in current Mamba-based methods with a random shuffle-based scanning method and design a new Shuffle Mamba framework. This approach mitigates the bias introduced by fixed scanning methods and provides an global receptive field for multi-modal image fusion. During testing, we employ Monte Carlo averaging to account for the introduced random factors. Extensive experiments on two tasks demonstrate that our approach outperforms state-of-the-art methods and exhibits strong generalization capabilities.

### v2

@@
-Shuffle Mamba: State Space Models with Random Shuffle for Multi-Modal Image Fusion
+arXiv:2409.01728v2[cs.CV] 27 Feb 2026
 Abstract
@@
 Although State-Space Models, such as Mamba, are proficient in long-range modeling with linear complexity, most Mamba-based approaches use fixed scanning strategies, which can introduce biased prior information.
-To mitigate this issue, we propose a novel Bayesian-inspired scanning strategy called Random Shuffle, supplemented by an theoretically-feasible inverse shuffle to maintain information coordination invariance, aiming to eliminate biases associated with fixed sequence scanning.
+To mitigate this issue, we propose a novel Bayesian-inspired scanning strategy called Random Shuffle, supplemented by a theoretically feasible inverse shuffle to maintain information coordination invariance, aiming to eliminate biases associated with fixed sequence scanning.
 Based on this transformation pair, we customized the Shuffle Mamba Framework, penetrating modality-aware information representation and cross-modality information interaction across spatial and channel axes to ensure robust interaction and an unbiased global receptive field for multi-modal image fusion.
 Furthermore, we develop a testing methodology based on Monte-Carlo averaging to ensure the model’s output aligns more closely with expected results.
-Extensive experiments across multiple multi-modal image fusion tasks demonstrate the effectiveness of our proposed method, yielding excellent fusion quality over state-of-the-art alternatives.
-Code will be available upon acceptance.
-Introduction
-Multi-modal image fusion, a fundamental task in computer vision, involves extracting and integrating valuable information from images of the same scene captured by different imaging modalities.
+Extensive experiments across multiple multi-modal image fusion tasks demonstrate the effectiveness of our proposed method, yielding excellent fusion quality compared to state-of-the-art alternatives.
+The code is available at https://github.com/caoke-963/Shuffle-Mamba.
+Index Terms:
+State-space model, Multi-modal Image fusion, Pan-sharpening
+I Introduction
+Multi-modal image fusion, a fundamental task in computer vision, involves extracting and integrating valuable information from images of the same scene captured by different imaging modalities [29, 27].
 This task aims to create a single composite image with a more comprehensive and informative representation, with typical applications including pan-sharpening and medical image fusion (MIF).
@@
 Specifically, PAN images offer superior spatial details but limited spectral resolution, while MS images provide abundant spectral resolution but lack spatial clarity.
-Through integrating the complementary information from both MS and PAN images into a composite representation, we can achieve an effective balance between spatial and spectral resolution.
+By integrating the complementary information from both MS and PAN images into a composite representation, we can achieve an effective balance between spatial and spectral resolution.
 Analogously, in the realm of MIF, various imaging technologies capture distinct types of information.
 For instance, Computed Tomography (CT) images deliver detailed insights into bones and high-density tissues, while Magnetic Resonance Imaging (MRI) offers higher-resolution images with rich soft tissue details.
-In virtue of this complementary information from various modalities, MIF can overcome the limitations of single-modality images, resulting in a more comprehensive and detailed representation for modern medical diagnosis.
-Figure 1: Visualization of Effective Receptive Fields (ERFs) for Conv, Self-Attention and our Method. A larger ERF is indicated by a more extensively distributed dark area.
+In virtue of this complementary information from various modalities, MIF can overcome the limitations of single-modality images, resulting in a more comprehensive and detailed representation of modern medical diagnosis.
+Figure 1: Visualization of Effective Receptive Fields (ERFs) and ERF Intensity Distributions for Conv, Self-Attention, Pan-Mamba, and Our Method.
+The upper part displays ERF visualizations, where larger, darker regions indicate broader receptive fields.
+The lower part presents the histogram of ERF intensities, with the x-axis representing ERF strength (lower values correspond to darker colors and stronger ERF responses) and the y-axis denoting pixel probability density.
+Additionally, the mean and standard deviation of ERF intensities are reported to provide a comprehensive comparison.
 In recent years, the prosperous advancement of deep neural networks (DNNs) has led to the development of numerous DNN-based multi-modal image fusion methods.
-In pan-sharpening, the pioneering work PNN (Masi et al.
-2016) employed a simple three-layer neural network to achieve remarkable results which was previously deemed unattainable, highlighting the superior learning capabilities of deep learning.
-After that,increasingly complex and deeper architectures have emerged (Zhou et al.
-2022b; Yang et al.
-2023), delivering excellent visual performance.
+In pan-sharpening, the pioneering work PNN [38] employed a simple three-layer neural network to achieve remarkable results that were previously deemed unattainable, highlighting the superior learning capabilities of deep learning.
+After that, increasingly complex and deeper architectures emerged [75, 66, 69], delivering excellent visual performance.
 Despite these advancements, existing multi-modal fusion methods face common limitations.
+Beyond improving fusion quality in standard scenarios, a growing trend in the field is to address image fusion in complex and adverse conditions.
+Recent studies have proposed unified frameworks to handle multiple complex scenes [24] and developed large-scale benchmarks to facilitate all-weather multi-modality fusion [26].
+Furthermore, incorporating global and local text perception has proven effective for fusion in adverse weather [25].
+These works highlight the increasing demand for robust perception capabilities in real-world applications.
 Convolutional neural network (CNN)-based approaches often struggle to establish global receptive fields.
@@
 CNNs are constrained by local receptive fields, which hinder their ability to model long-range dependencies.
-In contrast, ViT uses the self-attention mechanism to access a global receptive field but are burdened by quadratic computational complexity.
-Recently, structured state-space models have demonstrated enhanced capabilities in capturing long-term dependencies in sequence data while maintaining linear time complexity.
-Notably, Mamba has achieved significant strides in reducing inference latency and improving overall performance through selective state spaces and hardware-aware algorithms.
-With the introduction of vmamba (Liu et al.
-2024) and Vision Mamba (Zhu et al.
-2024), there has been a growing interest in applying state-space models to visual tasks.
-However, most current Mamba-based methods employ unidirectional SSMs, which endure certain limitations in their receptive field range due to their scanning approaches.
-Specifically, the receptive field is large for the initial portion of input patches but becomes significantly smaller for the latter portion, potentially compromising the model’s global modeling capabilities.
-Moreover, dissimilar to the dependencies between token orders in language modeling, the non-causal nature of 2-D spatial information in images presents a considerable challenge for simple sequential scanning methods.
-Traditional strategies, such as flattening image patches and scanning them sequentially, may introduce biased local 2D dependencies, thereby undermining the model’s ability to effectively identify spatial relationships.
-To address these challenges, we propose a new sequence scanning method called Random Shuffle Scanning.
-Figure 1 shows the visualization of the effective receptive field (ERF) of three methods.
-Compared to the Conv method, the Self-Attention method’s advantage lies in achieving a global receptive field.
-However, our method provides an overall unbiased global receptive field, enabling the network to discard fixed local priors and focus more effectively on what needs to be learned.
-For the sequential input image patches, we first apply position encoding, followed by a random shuffle of the patches before they are processed by the Mamba block for long-range dependency modeling.
-The random shuffle approach eliminates the preconception between local and global 2D dependencies in mathematical expectation, enabling the model to access an unbiased prior and ultimately establish a more consistent and global receptive field.
-Recognizing that shuffling the spatial position of the image patches may disrupt semantic information, we implement a corresponding inverse transformation to accurately restore the sequence order of the input patches after passing through the Mamba block.
-This information-preserving transformation pair underpins the key components of our model: Random Mamba Block, Random Channel Interactive Mamba Block, and Random Modal Interactive Mamba Block.
-Inspired by Dropout, we use Monte Carlo averaging to approximate the expected output, ensuring that the actual output at test time closely matches the anticipated results.
-Our contributions can be summarized as follows: 1) We design the Shuffle Mamba framework, where the random shuffle operation in the key component provides an expected unbiased global receptive field without increasing any parameters.
-2) We develop a specific strategy for training and testing this framework.
+In contrast, ViT uses the self-attention mechanism to access a global receptive field but is burdened by quadratic computational complexity.
+Structured state-space models have recently demonstrated enhanced capabilities in capturing long-term dependencies in sequence data while maintaining linear-time complexity.
+In particular, Mamba has made significant strides in reducing inference latency and improving overall performance through selective state spaces and hardware-aware algorithms.
+With the introduction of VMamba [33] and Vision Mamba [76], there has been a growing interest in applying state-space models to visual tasks.
+Most existing Mamba-based vision models adopt unidirectional or fixed scanning strategies to convert 2D spatial data into 1D sequences, which introduces fundamental limitations in receptive field modeling.
+Due to the sequential nature of state space models, earlier tokens typically benefit from broader receptive fields, while later tokens suffer from reduced context, leading to unbalanced global dependency modeling.
+Moreover, unlike the inherently causal structure of language, visual data is spatially non-causal and two-dimensional.
+Flattening image patches and scanning them sequentially disrupts spatial continuity, introduces orientation-specific biases, and hampers the modeling of local context.
+Prior studies, such as DgMamba [35] and EAMamb [30], have shown that fixed scanning breaks adjacent pixel relations, while MambaIR [10] and MS-VMamba [46] demonstrate that such strategies lead to uneven receptive field distributions and long-range information decay.
+These issues persist even with multi-directional scanning, indicating that deterministic scanning orders inherently limit the spatial modeling capacity of Mamba-based frameworks.
+To address these challenges, we propose a new sequence scanning method, termed Random Shuffle Scanning.
+Figure 1 visualizes the effective receptive fields (ERFs) of four representative approaches.
+Compared with convolutional models, self-attention achieves global perception but at the cost of quadratic computational complexity.
+Pan-Mamba achieves global modeling with linear complexity; however, its deterministic scanning order introduces structural bias, leading to an overemphasis on specific image patterns, such as horizontal stripes.
+In contrast, the proposed Shuffle Mamba establishes a balanced global receptive field with linear-time complexity, effectively removing fixed local priors and allowing the model to focus on task-relevant regions adaptively.
+To further investigate the receptive field distribution, the lower part of Figure 1 presents histograms of pixel-wise ERF intensities.
+The horizontal axis denotes gradient-based response intensity, while the vertical axis represents probability density.
+Compared with other methods, Shuffle Mamba exhibits a left-shifted distribution with a smaller mean value, corresponding to darker regions in the visualization and stronger activation responses.
+These results indicate that our randomized scanning promotes stronger global correlations and more uniform dependency modeling across spatial locations, effectively mitigating local bias and achieving theoretically unbiased global modeling.
+For the sequentially input image patches, we first apply position encoding and then randomly shuffle the patches before processing them through the Mamba block for long-range dependency modeling.
+This random shuffle operation removes the deterministic correlation between local and global 2D dependencies in expectation, enabling the model to access an unbiased prior and achieve consistent global perception.
+Recognizing that spatial shuffling may disrupt semantic consistency, a corresponding inverse transformation is applied to restore the original order of input patches after Mamba processing.
+This information-preserving shuffle–inverse pair serves as the foundation for the three key modules of our framework: the Random Mamba Block, Random Channel Interactive Mamba Block, and Random Modal Interactive Mamba Block.
+Furthermore, inspired by Dropout, we employ Monte Carlo averaging during inference to approximate the expected output, ensuring that the final prediction aligns closely with the theoretical expectation.
+Our contributions can be summarized as follows:
+  • We design the Shuffle Mamba framework, where the random shuffle operation in the key component provides an expected unbiased global receptive field without increasing any parameters.
+  • We develop a specific strategy for training and testing this framework.
 During training, each input is independently scanned using a random shuffle operation.
 During testing, we use Monte Carlo averaging to estimate the output of each Mamba block.
-3) Extensive experiments on two prominent multi-modal image fusion tasks demonstrate that our method accomplishes excellent performance in both quantitative assessments and visual quality.
-Related Works
-State Space Model
+  • Extensive experiments on two prominent multi-modal image fusion tasks demonstrate that our method accomplishes excellent performance in both quantitative assessments and visual quality.
+II Related Work
+II-A Pan-sharpening
+Traditional pan-sharpening methods can be categorized into three main groups: Component Substitution (CS)-based methods [19, 28, 7], Multi-Resolution Analysis (MRA)-based methods [44, 41], and Variational Optimization (VO)-based methods [5, 42].
+CS-based approaches extract the spatial details of the PAN image in the transform domain and inject them into the spatial components of the LRMS image.
+Conversely, MRA-based methods rely on a multi-resolution decomposition to derive the spatial information from the MS image and replace it with the content of the PAN image.
+VO-based models incorporate prior knowledge to formulate the fusion process and solve it iteratively.
+However, despite their diverse strategies, these traditional methods often experience significant spectral or spatial distortions due to inappropriate transformation schemes and ill-posed solution conditions.
+Inspired by SRCNN [4] in image super-resolution, PNN [38] adopted a simple model consisting of a three-layer convolutional neural network and achieved remarkable results, which has been seen as pioneering work for deep learning in pan-sharpening.
+This paved the way for many subsequent DL-based approaches with more complex structures.
+For instance, PanNet [67] conducts high-frequency filtering and residual connections to preserve detailed information and spectrum in the image.
+MSDCNN [68] incorporates a multi-scale and multi-depth network to extract various-level features.
+Meanwhile, SFINet [75] and MSDDN [14] leverage the Fourier transform and specially designed dual-domain structures to simultaneously capture spatial and frequency domain features, enabling deeper learning of high-frequency information.
+Transformer architecture has also been introduced to pan-sharpening [73], which employs self-attention mechanisms to enhance fusion quality.
+Additionally, considering the ill-posed property of the pan-sharpening problem, PanFlowNet [66] adopts a flow-based model to learn the conditional distribution of the fused image, which can generate multiple variations of HRMS images during inference.
+More recently, recognizing the limitations of previous methods in global information modeling and cross-modal information fusion, Pan-mamba [12] customizes the core components based on the state space model, yielding excellent results in feature extraction and fusion.
+II-B Medical Image Fusion
+Medical image fusion (MIF) plays a crucial role in improving disease diagnosis by combining essential information from different image sources into a single fused image.
+Since single-modality images capture limited information, MIF has become a vital tool to integrate diverse image features, facilitating more accurate and efficient diagnoses.
+Early methods for MIF predominantly relied on techniques such as spatial domain processing [45], transform domain approaches [22], and sparse representations [50].
+However, these traditional methods often struggled to capture deep semantic information and required manually designed fusion rules tailored to specific scenarios, which limited their generalizability.
+In recent years, deep learning techniques have gained increasing attention in the field of MIF.
+Liu et al.
+[32] were among the first to apply convolutional neural networks (CNNs) to MIF tasks, directly learning the mapping from source images to fused outputs.
+Building on this, Zero [20] utilized pre-trained CNNs to extract deep feature maps and adjust fusion weights by comparing these maps.
+To tackle the challenge of modeling global context in image fusion, MSRPAN [6] introduced a multi-scale pyramid network, residual connections, and attention mechanisms.
+This approach achieved superior visual quality and quantitative results by controlling the energy ratio during feature fusion.
+Furthermore, CDDFuse [71] introduced a correlation-driven Transformer-CNN dual-branch feature decomposition network, effectively separating modality-specific and modality-shared features.
+Inspired by the mixture of experts (MoE) model, TC-MoA [77] utilized task-specific routing strategies and customized adapters, achieving excellent performance across various image fusion tasks within a unified framework.
+Figure 2: The architecture of the proposed Shuffle Mamba framework.
+II-C State Space Model
 The State Space Model (SSM), originating from control theory, has found extensive applications in deep learning due to its remarkable ability to model long-range dependencies.
-Initially, the S4 (Gu, Goel, and Ré 2021) model introduced the concept of the SSM, effectively reducing the computational and memory requirements associated with state representation while enabling global information modeling.
-Building upon S4, the S5 (Smith, Warrington, and Linderman 2022) model features MIMO structure and efficient parallel scanning strategy to enhance performance without significantly increasing computational demands.
-Furthermore, the H3 (Mehta et al.
-2022) model further refines these methods, achieving competitive performance and efficiency comparable to the transformer in language modeling tasks.
-Figure 2: The architecture of the proposed Shuffle Mamba framework.
-Recently, Mamba (Gu and Dao 2023) has significantly improved inference speed and performance metrics through selective state spaces and hardware-aware algorithms.
-The introduction of Vmamba (Liu et al.
-2024) and Vision Mamba (Zhu et al.
-2024) has brought attention to the application of SSM in vision tasks.
-However, most existing SSM-based vision models (Zhu et al.
-2024; Ma, Li, and Wang 2024; Zheng and Wu 2024) employ a fixed scanning strategy, which may introduce preconceptions, particularly in low-level vision tasks (Xiao et al.
-2023).
-Specifically, this fixed order of choosing image patches can cause the model to gradually dismiss previous input sequences while processing the current patch, thereby compromising its capability to model global information.
-To alleviate this challenge, Vmamba (Liu et al.
-2024) introduced the CSM, which scans image pixels from various directions such as top-left, bottom-right, top-right, and bottom-left.
-Building on this idea, RSM (Zhao et al.
-2024) devised the OSSM to flatten image patches into sequences in eight directions, enhancing the network’s capability to capture and model large-scale spatial features.
-Additionally, LocalMamba (Huang et al.
-2024) applies the windowed selective scan to ensure a harmonious integration of global and local visual cues.
-RSMamba (Chen et al.
-2024) incorporates dynamic multi-path activation mechanisms to model non-causal data.
+Initially, the S4 [9] model introduced the concept of the SSM, effectively reducing the computational and memory requirements associated with state representation while enabling global information modeling.
+Building upon S4, the S5 [47] model features a MIMO structure and an efficient parallel scanning strategy to enhance performance without significantly increasing computational demands.
+Furthermore, the H3 [39] model further refines these methods, achieving competitive performance and efficiency comparable to the transformer in language modeling tasks.
+Recently, Mamba [8] has significantly improved inference speed and performance metrics through selective state spaces and hardware-aware algorithms.
+The introduction of Vmamba [33] and Vision Mamba [76] has brought attention to the application of SSM in vision tasks.
+However, most existing SSM-based vision models [76, 37, 72] employ a fixed scanning strategy, which may introduce preconceptions, particularly in low-level vision tasks [57].
+Specifically, this fixed order of selecting image patches can cause the model to gradually disregard previous input sequences while processing the current patch, thereby compromising its ability to model global information.
+To alleviate this challenge, Vmamba [33] introduced the CSM, which scans image pixels from various directions such as top-left, bottom-right, top-right, and bottom-left.
+Building on this idea, RSM [70] introduced the OSSM to transform image patches into sequences in eight directions, thereby enhancing the network’s ability to capture and model large-scale spatial features.
+Additionally, LocalMamba [16] applies the windowed selective scan to ensure a harmonious integration of global and local visual cues.
+RSMamba [3] incorporates dynamic multi-path activation mechanisms to model non-causal data.
 However, these methods do not simultaneously consider the integrity of the image’s structure and the randomness of the pixels during scanning processing, and are essentially still special fixed strategies.
-To address this limitation, we introduced the Shuffle Mamba framework, which includes random shuffle and inverse operations to obtain a global receptive field without bias in expectations.
-Method
-Preliminaries
+To address this limitation, we introduced the Shuffle Mamba framework, which employs random shuffling and inverse operations to achieve a global receptive field without introducing expected biases.
+As a result, our approach not only advances pansharpening and medical image fusion but also demonstrates significant potential for a wider range of low-level vision tasks, including infrared and visible image fusion [2, 1, 17], pan-denoising [63, 64], and multi-modal UAV image fusion [18].
+III Methods
+In this section, we introduce the foundational concepts of state-space models.
+Next, we describe the training strategy and module design underlying the proposed random shuffle framework.
+Finally, we detail the Monte Carlo averaging approach used during testing and discuss the design of the loss function tailored to different tasks.
+III-A Preliminaries
 Inspired by continuous linear time-invariant (LTI) systems, SSMs exploit an implicit latent state $h(t)\in{\mathbf{R}^{N}}$to map a 1-D sequence $x(t)\in\mathbf{R}$ to $y(t)\in\mathbf{R}$.
@@
 where $\mathbf{L}$ represents the length of the sequence $x$, $\bar{\mathbf{K}}\in\mathbf{{R}^{L}}$ is a structured convolution kernel.
-Network Architecture
+Figure 3: The Random Shuffle Scanning for training.
+Figure 4: The Monte-Carlo averaging for testing.
+III-B Network Architecture
 The proposed Shuffle Mamba framework consists of three functional components: the Random Mamba block (RM block), the Random Channel Interactive Mamba block (RCIM block), and the Random Modal Interactive Mamba block (RMIM block).
@@
 Given the limited receptive field of the convolutional layers, which makes capturing global features challenging, we perform patch embedding and send the resulting patches to the RM block for global feature extraction.
-This process yields the global modality-specific features $\mathbf{F}_{n}^{m}$ and $\mathbf{F}_{n}^{p}$, where $PE=PatchEmbed$:
-$\mathbf{F}_{0}^{m},\mathbf{F}_{0}^{p}=PE(\phi({\mathbf{M}_{up}})),PE(\phi({\mathbf{P}_{in}}))$ (8)
-$\mathbf{F}_{n}^{m}={{\psi}_{1n}}\cdot\cdot\cdot({{\psi}_{11}}({{\psi}_{10}}(\mathbf{F}_{0}^{m}))$ (9)
-$\mathbf{F}_{n}^{p}={{\psi}_{2n}}\cdot\cdot\cdot({{\psi}_{21}}({{\psi}_{20}}(\mathbf{F}_{0}^{p}))$ (10)
+This process yields the global modality-specific features $\mathbf{F}_{n}^{m}$ and $\mathbf{F}_{n}^{p}$, where $PE$ means the process of $PatchEmbed$, $\phi$ denotes the initial convolution operation, $\psi$ corresponds to the Random Mamba module, and the subscripts indicate two branches and their sequential modules.
+$\mathbf{F}_{0}^{m},\mathbf{F}_{0}^{p}=PE(\phi({\mathbf{M}_{up}})),PE(\phi({\mathbf{P}_{in}})),$ (8)
+$\mathbf{F}_{n}^{m}={{\psi}_{1n}}\cdot\cdot\cdot({{\psi}_{11}}({{\psi}_{10}}(\mathbf{F}_{0}^{m})),$ (9)
+$\mathbf{F}_{n}^{p}={{\psi}_{2n}}\cdot\cdot\cdot({{\psi}_{21}}({{\psi}_{20}}(\mathbf{F}_{0}^{p})).$ (10)
 The global modality-specific features are then sent to the RCIM block for simple channel information exchange, without introducing additional parameters.
@@
 Thus, the final fused image $\mathbf{H}_{out}$} can be accessed by reshaping and residual connection:
-$\mathbf{F}_{k+j}^{m}={{\theta}_{j}}\cdot\cdot\cdot({{\theta}_{1}}({{\theta}_{0}}(\mathbf{F}_{k}^{m},\mathbf{F}_{k}^{p}),\mathbf{F}_{k}^{p}),\mathbf{F}_{k}^{p})$ (11)
-${\mathbf{H}_{out}}=\phi(reshape(\mathbf{F}_{k+j}^{m}))+{\mathbf{M}_{up}}$ (12)
-Figure 3: The Random Shuffle Scanning for training.
-Figure 4: The Monte-Carlo averaging for testing.
-Key Components
-Random Shuffle Scanning.
+$\mathbf{F}_{k+j}^{m}={{\theta}_{j}}\cdot\cdot\cdot({{\theta}_{1}}({{\theta}_{0}}(\mathbf{F}_{k}^{m},\mathbf{F}_{k}^{p}),\mathbf{F}_{k}^{p}),\mathbf{F}_{k}^{p}),$ (11)
+${\mathbf{H}_{out}}=\phi(reshape(\mathbf{F}_{k+j}^{m}))+{\mathbf{M}_{up}}.$ (12)
+III-C Key Components
+III-C1 Random Shuffle Scanning
 Mamba was originally designed to adapt to the modeling of language sequences.
-We devised a Random Shuffle Scanning approach to access unbiased local and global dependencies in 2D image while ensuring a global receptive field similar to self-attention.
+We devised a Random Shuffle Scanning approach to access unbiased local and global dependencies in the 2D image while ensuring a global receptive field similar to self-attention.
 As shown in Figure 3, for sequentially input 2D image patches, we first apply depth-wise convolution for position modeling.
@@
 For this reason, the random shuffle and its inverse operation constitute an information-lossless transformation pair.
-Random Mamba Block.
+III-C2 Random Mamba Block
 Based on this shuffle-inverse pair, we designed the Random Mamba Block.
@@
 Finally, we apply the inverse shuffle and residual connection to get the final output sequence $\mathbf{F}_{l}$}.
-Random Channel Interactive Mamba Block.
-In the RCIM block, we refer to the approach from (He et al.
-2024a) to achieve lightweight feature interaction between different modalities.
+III-C3 Random Channel Interactive Mamba Block
+In the RCIM block, we refer to the approach from [12] to achieve lightweight feature interaction between different modalities.
 We use the split operation to divide modality features $\mathbf{F}_{k}^{m}$ and $\mathbf{F}_{k}^{p}$ into two halves based on the channel dimension, followed by complementary splicing.
@@
 By repeating these steps, the global modality-specific features are initially fused.
-Random Modal Interactive Mamba block.
+III-C4 Random Modal Interactive Mamba Block
 Motivated by cross-attention, we designed the RMIM block for processing multi-modal image information.
@@
 Finally, the output $\mathbf{F}_{l}^{m}$ of the module can be obtained through depth-wise convolution and feature flattening.
-Testing with Monte Carlo averaging
-Table 1: Quantitative comparison of pan-sharpening task on three datasets.
-Bold and underline show the best and second-best values, respectively.
+III-D Testing with Monte Carlo averaging
+We incorporate stochastic factors into the random shuffle operation, necessitating the marginalization of these factors in the generation of the final fusion.
+However, the random shuffle method presents a theoretical challenge due to the exponentially large number of potential models, making precise averaging of their predictions infeasible.
+Drawing inspiration from dropouts [48], we approximate the expected value of the entire model through layered expectations.
+Therefore, the computation of the random shuffle during testing can be expressed as follows, where $\mathbf{RM}=RandomMamba$:
+$\mathbf{R}{\mathbf{M}^{test}}(x)=\underset{\mathbf{\mathcal{S}}}{\mathop{\mathbb{E}}}\,[\mathbf{RM}(x,\mathbf{\mathcal{S}})]$ (13)
+Estimating $\mathbf{R}{{\mathbf{M}}^{test}}$ based on the aforementioned equation requires enumerating all possible shuffle results, which imposes a significant computational burden.
+Therefore, we employ Monte Carlo averaging to estimate its expectation:
+$\mathbf{R}{\mathbf{M}^{test}}(x)\approx\frac{1}{M}\sum\limits_{i=1}^{M}{\mathbf{RM}(x,{{\mathbf{\mathcal{S}}}_{i}})}$ (14)
+Specifically, the input image is independently shuffled M{M} times, and then the M{M} outputs of $\mathbf{RM}$ are calculated.
+The mean of these outputs is computed to obtain the final estimate.
+When ${M}\to\infty$ is present, the Monte Carlo estimator closely approximates the true mean.
+Figure 4 illustrates the test process we designed, which significantly reduces the actual testing time by incorporating multiple identical inputs in a mini-batch and utilizing GPUs for parallel computation.
+TABLE I: Quantitative comparison of pan-sharpening task on three datasets.
+The best results are highlighted in bold.
+The second-best results are highlighted in underline.
 $\uparrow$ indicates that the larger the value, the better the performance, and $\downarrow$ indicates that the smaller the value, the better the performance.
@@
 […]
-INNformer AAAI’22 41.6903 0.9704 0.0227 0.9514 47.3528 0.9893 0.0102 0.5479 30.5365 0.9225 0.0747 3.1142
-SFINet ECCV’22 41.7244 0.9725 0.0220 0.9506 47.4712 0.9901 0.0102 0.5479 30.5901 0.9236 0.0741 3.0798
-MSDDN TGRS’23 41.8435 0.9711 0.0222 0.9478 47.4101 0.9895 0.0101 0.5414 30.8645 0.9258 0.0757 2.9581
-PanFlowNet ICCV’23 41.8548 0.9712 0.0224 0.9335 47.2533 0.9884 0.0103 0.5512 30.4873 0.9221 0.0751 2.9531
-FAME AAAI’24 42.0262 0.9723 0.0215 0.9172 47.6721 0.9898 0.0098 0.5542 30.9903 0.9287 0.0697 2.9531
-DISPNet AAAI’24 41.8768 0.9702 0.0221 0.9157 47.4529 0.9898 0.0111 0.5532 30.0426 0.9153 0.0776 3.2620
-Pan-mamba arxiv’24 42.2354 0.9729 0.0212 0.8975 47.6453 0.9894 0.0103 0.5286 31.1740 0.9302 0.0698 2.8910
-Ours - 42.3428 0.9734 0.0208 0.8840 47.9180 0.9903 0.0097 0.5109 31.4005 0.9327 0.0676 2.8098
-Table 2: Evaluation of our method on real-world full-resolution scenes from the WV2 dataset. Bold and underline show the best and second-best values, respectively.
-Metric SFIM Brovey GFPCA INNformer SFINet PanFlowNet FAME DISPNet Pan-mamba Ours
-$D_{\lambda}$downarrow 0.1403 0.1026 0.1139 0.0995 0.1034 0.0966 0.0951 0.0944 0.0966 0.0941
-$D_{S}$downarrow 0.1320 0.1409 0.1535 0.1305 0.1305 0.1274 0.1263 0.1264 0.1272 0.1266
-QNR $\uparrow$ 0.7826 0.7728 0.7532 0.7858 0.7827 0.7910 0.7933 0.7938 0.7911 0.7939
+SFIM 34.1297 0.8975 0.0439 2.3449 36.9060 0.8882 0.0318 1.7398 21.8212 0.5457 0.1208 8.9730
+Brovey 35.8646 0.9216 0.0403 1.8238 37.2260 0.9034 0.0309 1.6736 22.5060 0.5466 0.1159 8.2331
+IHS 35.6376 0.9176 0.0423 1.8774 37.7974 0.9026 0.0218 1.3720 22.5608 0.5470 0.1217 8.2433
+GS 35.2962 0.9027 0.0461 2.0278 38.1754 0.9100 0.0243 1.5336 22.5579 0.5354 0.1266 8.3616
+GFPCA 34.5581 0.9038 0.0488 2.1411 37.9443 0.9204 0.0314 1.5604 22.3344 0.4826 0.1294 8.3964
+SFNLR 33.8636 0.8957 0.0481 2.2415 36.2367 0.8850 0.0363 1.8611 21.7647 0.5636 0.1494 9.0962
+LRTCFPan 34.7756 0.9112 0.0426 2.0075 36.9253 0.8946 0.0332 1.7060 22.1574 0.5735 0.1380 8.6796
+MSDCNN 41.3355 0.9664 0.0242 0.9940 45.6847 0.9827 0.0135 0.6389 30.3038 0.9184 0.0782 3.1884
+INNformer 41.6903 0.9704 0.0227 0.9514 47.3528 0.9893 0.0102 0.5479 30.5365 0.9225 0.0747 3.1142
+SFINet 41.7244 0.9725 0.0220 0.9506 47.4712 0.9901 0.0102 0.5479 30.5901 0.9236 0.0741 3.0798
+MSDDN 41.8435 0.9711 0.0222 0.9478 47.4101 0.9895 0.0101 0.5414 30.8645 0.9258 0.0757 2.9581
+PanFlowNet 41.8548 0.9712 0.0224 0.9335 47.2533 0.9884 0.0103 0.5512 30.4873 0.9221 0.0751 2.9531
+FAME 42.0262 0.9723 0.0215 0.9172 47.6721 0.9898 0.0098 0.5542 30.9903 0.9287 0.0697 2.9531
+DISPNet 41.8768 0.9702 0.0221 0.9157 47.4529 0.9898 0.0111 0.5532 30.0426 0.9153 0.0776 3.2620
+Pan-mamba 42.2354 0.9729 0.0212 0.8975 47.6453 0.9894 0.0103 0.5286 31.1740 0.9302 0.0698 2.8910
+ARConv 41.8281 0.9704 0.0222 0.9395 47.3699 0.9887 0.0102 0.5492 30.8156 0.9251 0.0717 3.0114
+Ours 42.3428 0.9734 0.0208 0.8840 47.9180 0.9903 0.0097 0.5109 31.4005 0.9327 0.0676 2.8098
+III-E Loss Function
+In accordance with established norms within this area, the loss function used in our model for pan-sharpening is the L1 loss.
+Assume the two input images of the model be represented as $\mathbf{M}_{in}$} and $\mathbf{P}_{in}$}, with $\mathbf{M}_{up}$} obtained by upsampling $\mathbf{M}_{in}$}.
+The model output is denoted as $\mathbf{H}_{out}$}, and the corresponding ground truth or basic reference features (if available) are represented as $\mathbf{H}_{gt}$}, the loss of pan-sharpening $\mathbf{L}_{pan}$} can be articulated as follows:
+$\mathbf{\mathcal{L}_{pan}}={{\left\|{\mathbf{H}_{gt}}-{\mathbf{H}_{out}}\right\|}_{1}}$ (15)
+In the MIF task, we utilize two input images and the fused image to compute a composite loss function $\mathbf{L}_{fus}$} comprising intensity loss $\mathbf{L}_{1}$}, structural similarity index (SSIM) loss $\mathbf{L}_{ssim}$}, and gradient loss $\mathbf{L}_{grad}$}.
+${\mathbf{L}_{1}}={{\left\|{\mathbf{H}_{out}}-\max({\mathbf{M}_{up}},{\mathbf{P}_{in}})\right\|}_{1}}$ (16)
+${\mathbf{L}_{ssim}}=0.5\times\mathbf{s}({\mathbf{H}_{out}},{{M}_{up}})+0.5\times\mathbf{s}({\mathbf{H}_{out}},{{P}_{in}})$ (17)
+$gra{{d}_{join}}=\max(\mathbf{sobel}({\mathbf{M}_{up}}),\mathbf{sobel}({\mathbf{P}_{in}}))$ (18)
+${\mathbf{L}_{grad}}={{\left\|\mathbf{sobel}({\mathbf{H}_{out}})-gra{{d}_{join}}\right\|}_{1}}$ (19)
+${\mathbf{L}_{fus}}=20{\mathbf{L}_{1}}+50{\mathbf{L}_{ssim}}+100{\mathbf{L}_{grad}}$ (20)
+Here, $\mathbf{s}$ denotes the structural similarity index (SSIM) calculation operation on the image, while $\mathbf{sobel}$ represents the Sobel convolution operation.
+TABLE II: Evaluation of our method on real-world full-resolution scenes from the WV2 dataset. Bold and underline show the best and second-best values, respectively.
+Metric SFIM Brovey GFPCA INNformer SFINet PanFlowNet FAME DISPNet Pan-mamba ARConv Ours
+$D_{\lambda}$ $\downarrow$ 0.1403 0.1026 0.1139 0.0995 0.1034 0.0966 0.0951 0.0944 0.0966 0.0963 0.0941
+$D_{S}$ $\downarrow$ 0.1320 0.1409 0.1535 0.1305 0.1305 0.1274 0.1263 0.1264 0.1272 0.1276 0.1266
+QNR $\uparrow$ 0.7826 0.7728 0.7532 0.7858 0.7827 0.7910 0.7933 0.7938 0.7911 0.7910 0.7939
 Figure 5: Comparative visual experiments of several methods on WV3 datasets
-We incorporate stochastic factors into the random shuffle operation, necessitating the marginalization of these factors in the generation of the final fusion.
-However, the random shuffle method presents a theoretical challenge due to the exponentially large number of potential models, making precise averaging of their predictions infeasible.
-Drawing inspiration from dropouts (Srivastava et al.
-2014), we approximate the expected value of the entire model through layered expectations.
-Therefore, the computation of the random shuffle during testing can be expressed as follows, where $\mathbf{RM}=RandomMamba$:
-$\mathbf{R}{\mathbf{M}^{test}}(x)=\underset{\mathbf{\mathcal{S}}}{\mathop{\mathbb{missing}}{E}}\,[\mathbf{RM}(x,\mathbf{\mathcal{S}})]$ (13)
-In fact, estimating $\mathbf{R}{{\mathbf{M}}^{test}}$ based on the aforementioned equation requires enumerating all possible shuffle results, which imposes a significant computational burden.
-Therefore, we employ Monte Carlo averaging to estimate its expectation:
-$\mathbf{R}{\mathbf{M}^{test}}(x)\approx\frac{1}{M}\sum\limits_{i=1}^{M}{\mathbf{RM}(x,{{\mathbf{\mathcal{S}}}_{i}})}$ (14)
-Specifically, the input image is independently shuffled ${M}$ times, and then the ${M}$ outputs of $\mathbf{RM}$ are calculated.
-The mean of these outputs is computed to obtain the final estimate.
-When ${M}\to\infty$ is present, the Monte Carlo estimator closely approximates the true mean.
-Figure 4 illustrates the testing process we designed, which significantly reduces the actual testing time by incorporating multiple identical inputs in a mini-batch and utilizing GPUs for parallel computation.
-Table 3: Quantitative comparison of MIF task on three datasets. bold and underline show the best and second-best values.
+Figure 6: Comparative visual experiments of several methods on WV2 datasets
+Figure 7: Comparative visual experiments of several methods on GF2 datasets
+Figure 8: The result of our approach on full-resolution WV2 dataset.
+IV Experiments
+IV-A Datasets and Benchmark
+For the pan-sharpening task, we utilized datasets from WorldView-II (WV2), GaoFen2 (GF2), and WorldView-III (WV3), encompassing a diverse range of urban and natural scenes.
+Training samples were prepared using the Wald protocol [51], which is widely adopted when the ground truth is inaccessible.
+To evaluate the effectiveness of our proposed method, we conducted a comprehensive comparison against both traditional and deep learning-based approaches.
+The baseline methods included classical algorithms such as SFIM [34], Brovey [7], GS [19], IHS [11], GFPCA [28], SFNLR [58] and LRTCFPan [56],as well as state-of-the-art deep learning techniques, including MSDCNN [68], INNformer [74], SFINet [75], MSDDN [14], PanFlowNet [66], FAME [13], DISPNet [52], Pan-Mamba [12], and ARConv [54].
+TABLE III: Evaluation of parameter count and GFLOPs.
+Methods SRPPNN INNformer MSDDN PanFlowNet FAME DISPNet Panmamba ARConv Ours
+Params(M) 1.7114 0.0706 0.3185 0.0873 0.6286 0.4433 0.1827 4.4147 0.2077
+GFLOPs(G) 21.1059 1.3079 2.5085 5.7126 10.1943 7.7659 3.0088 14.0833 3.4021
+Infer Time(ms) 6.7071 11.9172 7.4857 5.3810 12.7889 11.5221 19.0202 63.3764 24.4713
+TABLE IV: Quantitative comparison of MIF task on three datasets. bold and underline show the best and second-best values.
 PET CT SPECT
 […]
-PSLPT 0.888 0.548 0.373 0.815 0.675 0.502 0.432 0.810 0.850 0.359 0.325 0.933
-EMFusion 0.943 0.685 0.783 1.221 1.190 0.552 0.475 1.266 0.885 0.665 0.692 1.212
-MSRPAN 1.017 0.581 0.799 1.182 1.319 0.436 0.455 1.261 0.960 0.525 0.560 1.153
-SwinFusion 1.642 0.703 0.683 0.725 1.537 0.522 0.545 0.579 1.678 0.744 0.720 0.684
-Zero 0.950 0.635 0.774 1.162 0.768 0.320 0.582 1.199 1.046 0.582 0.681 1.180
-U2Fusion 0.947 0.460 0.292 0.494 0.309 0.074 0.489 0.042 0.865 0.419 0.696 0.479
-CDDFuse 1.481 0.650 0.765 1.227 1.589 0.526 0.530 1.224 0.995 0.786 0.719 1.169
-Ours 1.491 0.797 0.741 1.256 1.580 0.546 0.592 1.330 1.470 0.820 0.747 1.240
+EMFusion 0.943 0.685 0.783 0.611 1.190 0.552 0.475 0.633 0.885 0.665 0.692 0.606
+MSRPAN 1.017 0.581 0.799 0.591 1.319 0.436 0.455 0.631 0.960 0.525 0.560 0.577
+SwinFusion 1.642 0.703 0.683 0.363 1.537 0.522 0.545 0.290 1.678 0.744 0.720 0.342
+Zero 0.950 0.635 0.774 0.581 0.768 0.320 0.582 0.600 1.046 0.582 0.681 0.590
+U2Fusion 0.947 0.460 0.292 0.247 0.309 0.074 0.489 0.021 0.865 0.419 0.696 0.240
+CDDFuse 1.481 0.650 0.765 0.614 1.589 0.526 0.530 0.612 0.995 0.786 0.719 0.585
+PSLPT 0.888 0.548 0.373 0.408 0.675 0.502 0.432 0.405 0.850 0.359 0.325 0.467
+Fusionmamba 1.399 0.791 0.736 0.618 1.532 0.538 0.587 0.664 1.368 0.761 0.732 0.619
+TC-MoA 1.387 0.540 0.615 0.563 1.584 0.451 0.503 0.414 1.487 0.554 0.594 0.557
+PRRGAN 1.405 0.641 0.696 0.607 1.477 0.502 0.497 0.563 1.109 0.682 0.701 0.589
+TFS-Diff 1.489 0.665 0.758 0.621 1.591 0.530 0.547 0.658 1.549 0.743 0.684 0.611
+Ours 1.491 0.797 0.741 0.628 1.580 0.546 0.592 0.665 1.470 0.820 0.747 0.620
+TABLE V: Quantitative comparison of IVIF task on three datasets. bold and underline show the best and second-best values.
+MSRS RoadScene M3FD
+[…]
+SemLA 1.392 0.608 0.290 0.498 1.269 0.564 0.415 0.518 1.495 0.542 0.363 0.473
+U2Fusion 1.258 0.512 0.391 0.440 1.605 0.564 0.506 0.546 1.753 0.673 0.578 0.463
+SwinFusion 1.647 0.825 0.558 0.504 1.576 0.614 0.450 0.534 1.588 0.746 0.616 0.492
+CDDFuse 1.549 0.819 0.548 0.459 1.707 0.610 0.450 0.515 1.673 0.802 0.613 0.460
+PSLPT 1.374 0.753 0.553 0.501 1.009 0.134 0.171 0.238 0.638 0.958 0.321 0.483
+Fusionmamba 1.635 0.974 0.652 0.511 1.322 0.635 0.543 0.519 1.414 0.747 0.580 0.480
+TC-MoA 1.661 0.811 0.565 0.515 1.562 0.577 0.477 0.522 1.556 0.579 0.508 0.466
+MLFuse 1.520 0.753 0.519 0.498 1.595 0.629 0.527 0.545 1.600 0.592 0.460 0.501
+Up-Fusion 1.584 0.813 0.602 0.502 1.718 0.644 0.536 0.533 1.616 0.816 0.621 0.496
+Ours 1.663 0.983 0.655 0.526 1.708 0.651 0.540 0.548 1.766 0.952 0.619 0.515
+In the MIF task, we employed medical images from the Harvard Medical website, comprising pairs such as MRI-CT, MRI-PET, and MRI-SPECT [40, 55].
+For this task, we compared our method with various deep learning-based multi-modal fusion techniques, including EMFusion [62], MSRPAN [6], SwinFusion [36], Zero [20], U2Fusion [60], CDDFuse [71], PSLPT [53], Fusionmamba [43], TC-MoA [77], PRRGAN [15], and TFS-Diff [65].
 Figure 9: Comparative visual experiments of several methods on MRI-PET datasets
-Table 4: Ablation for our methods on three datasets. The PSNR/SSIM/SAM/ERGAS values on benchmarks are reported. The best results are shown in bold.
+Figure 10: Comparative visual experiments of several methods on MRI-SPECT datasets
+IV-B Implement Details
+All experiments were conducted using the PyTorch framework on two NVIDIA RTX 4090 GPUs.
+We trained for 500 epochs with a batch size of 4 for the pan-sharpening task and 200 epochs with a batch size of 1 for the medical image fusion (MIF) task.
+The network parameters were optimized using the Adam optimizer.
+The initial learning rate was set to 5e45e-4, which was subsequently reduced to 5e85e-8 using the CosineAnnealingLR scheduler over the specified epochs.
+For the pan-sharpening task, we randomly cropped training set images to obtain LRMS patches of size 32x32 and PAN images of size 128x128.
+For the MIF task, the training set images were cropped to 256x256.
+Figure 11: Comparative visual experiments of several methods on MRI-CT datasets
+TABLE VI: Ablation for our methods on three datasets. The PSNR/SSIM/SAM/ERGAS values on benchmarks are reported. The best results are shown in bold. RSS denotes Random-Shuffle Scanning.
 Ablation Variant WorldView-II Gaofen-2 WorldView-III
 Baseline - 42.3428/0.9734/0.0208/0.8840 47.9180/0.9903/0.0097/0.5109 31.4005/ 0.9327/0.0676/2.8098
-Core Operation RM w/o Random Shuffle 42.1460/0.9723/0.0211/0.9042 47.4284/0.9890/0.0102/0.5355 31.3915/0.9329/0.0676/2.8113
-RCIM w/o Random Shuffle 42.2443/0.9729/0.0210/0.8919 47.6847/0.9899/0.0099/0.5292 31.3476/0.9325/0.0680/2.8230
-RMIM w/o Random Shuffle 42.2136/0.9727/0.0211/0.8978 47.5787/0.9895/0.0100/0.5324 31.0971/0.9295/0.0704/2.9097
-Loss Function
-In accordance with established norms within this area, the loss function employed in our model for pan-sharpening is the L1 loss.
-In the MIF task, we use two input images and the fused image to calculate a composite loss function that includes L1 loss, SSIM loss, and gradient loss.
-Experiments
-Datasets and Benchmark
-For the pan-sharpening task, we utilized datasets from WorldView-II (WV2), GaoFen2 (GF2), and WorldView-III (WV3), encompassing a variety of urban and natural scenes.
-We generated training samples following the Wald (Wald, Ranchin, and Mangolini 1997) protocol in the inaccessible ground truth.
-We conducted a thorough comparison of our proposed method against classical approaches and deep learning-based methods, including INNformer (Zhou et al.
-2022a), SFINet (Zhou et al.
-2022b), MSDDN (He et al.
-2023), PanFlowNet (Yang et al.
-2023), FAME (He et al.
-2024b), DISPNet (Wang et al.
-2024) and Pan-mamba (He et al.
-2024a).
-In the MIF task, we employed medical images from the Harvard Medical website, comprising pairs such as MRI-CT, MRI-PET, and MRI-SPECT.
-For this task, we compared our method with various deep learning-based multi-modal fusion techniques, including PSLPT (Wang, Deng, and Vivone 2024), EMFusion (Xu and Ma 2021), MSRPAN (Fu et al.
-2021), SwinFusion (Ma et al.
-2022), Zero (Lahoud and Süsstrunk 2019), U2Fusion (Xu et al.
-2020), and CDDFuse (Zhao et al.
-2023).
-Implementation Details
-All experiments were conducted using the PyTorch framework on two NVIDIA RTX 3060 GPUs.
-We trained for 500 epochs with a batch size of 4 for the pan-sharpening task and 200 epochs with a batch size of 1 for the medical image fusion (MIF) task.
-The network parameters were optimized using the Adam optimizer.
-The initial learning rate was set to $45e-$ 4, which was subsequently reduced to $85e-$ 8 using the CosineAnnealingLR scheduler over the specified epochs.
-For the pan-sharpening task, we randomly cropped training set images to obtain LRMS patches of size 32x32 and PAN images of size 128x128.
-For the MIF task, the training set images were cropped to 256x256.
-Comparison with SOTA Methods
-Pan-sharpening.
-The experimental results on three datasets are presented in Table 1.
-Reference metrics, including PSNR, SSIM, SAM, and ERGAS (Alparone et al.
-2007), are used to evaluate the fusion effect.
+Single Operation (1) RM w/o RSS 42.1460/0.9723/0.0211/0.9042 47.4284/0.9890/0.0102/0.5355 31.3915/0.9329/0.0676/2.8113
+(2) RCIM w/o RSS 42.2443/0.9729/0.0210/0.8919 47.6847/0.9899/0.0099/0.5292 31.3476/0.9325/0.0680/2.8230
+(3) RMIM w/o RSS 42.2136/0.9727/0.0211/0.8978 47.5787/0.9895/0.0100/0.5324 31.0971/0.9295/0.0704/2.9097
+Multi Operations (1) + (2) 42.1205/0.9719/0.0214/0.9075 47.4000/0.9878/0.0104/0.5405 31.2150/0.9316/0.0692/2.8782
+(1) + (3) 42.1128/0.9719/0.0215/0.9086 47.3851/0.9876/0.0105/0.5420 31.0547/0.9292/0.0706/2.9053
+(2) + (3) 42.1684/0.9721/0.0213/0.9037 47.4655/0.9889/0.0103/0.5381 31.1200/0.9294/0.0702/2.9019
+(1) + (2) + (3) 42.0765/0.9718/0.0218/0.9094 47.3111/0.9857/0.0107/0.5498 31.0026/0.9274/0.0719/2.9142
+TABLE VII: Ablation of different scanning strategies.
+The best results are shown in bold.
+RSS denotes Random-Shuffle Scanning, $S$ denotes Sequential Scanning, BD denotes Bidirectional Scanning, and DS denotes Diagonal Scanning
+WorldView-II
+[…]
+RSS -> SS 42.0967 0.9718 0.0218 0.9094
+RSS -> BS 42.2294 0.9726 0.0211 0.8933
+RSS -> DS 42.1539 0.9721 0.0214 0.8967
+Ours (RSS) 42.3428 0.9734 0.0208 0.8840
+IV-C Comparison with SOTA Methods
+IV-C1 Pan-sharpening
+The experimental results on three datasets are presented in Table I.
+Reference metrics, including PSNR, SSIM, SAM, and ERGAS, are used to evaluate the fusion effect.
 The results demonstrate that the proposed method outperforms the SOTA methods across all metrics.
-Notably, in PSNR metric, our method achieves improvements of 0.1047, 0.2727, and 0.2301 dB over Pan-mamba, which has the second-best performance.
-Better PSNR and SSIM evidence that the fusion results transfers more information from the original image and experiences less distortion.
-In the qualitative comparison, Figure 5 shows a representative sample from the WV3 dataset.
-Our method performs better on the MSE graph, indicating that the fusion result is closer to the ground truth.
-Compared to other methods, our approach achieves a more accurate restoration of spectral and spatial details, highlighting the advantages of our fusion technique.
-To further verify the generalization ability of our method in full-resolution scenes, we use three non-reference metrics, including ${{D}_{s}}$, ${{D}_{\lambda}}$ and QNR, to evaluate our method on the full WorldView-II dataset.
-Table 2 presents the experimental results.
-Our method outperforms the comparison methods across all results, demonstrating the strong adaptability of Shuffle Mamba in image fusion.
-Medical Image Fusion.
-The quantitative comparison results of four metrics on the MIF dataset are shown in Table 3.
-The proposed method performs well across almost all metrics, demonstrating its effectiveness in medical image fusion.
-In our experimental results, a higher VIF indicates closer alignment with human perception.
-Improved SCD, Qabf, and SSIM scores suggest that the fused image maintains higher similarity and experiences less distortion compared to the original images.
-Figure 6 presents a qualitative comparison of several methods on the MRI-PET dataset.
-Our method exhibits superior visual quality, a finding that is corroborated by the experimental metrics.
-Ablation Experiments
-To verify the effectiveness of the Random Shuffle operation, we conducted corresponding ablation experiments.
-The core operations of these experiments involved removing the shuffle operations from the three main modules.
-Notably, in the experiment where RM-related components were removed, the RM block in the RCIM module was excluded.
-The experimental results are presented in Table 4.
-Removing the shuffle operation resulted in a significant decrease in model performance, indicating that the new scanning strategy designed under the Shuffle Mamba framework effectively enhances the quality of multi-modal image fusion.
+Notably, in the PSNR metric, our method achieves improvements of 0.1047, 0.2727, and 0.2301 dB over Pan-mamba, which has the second-best performance.
+Better PSNR and SSIM evidence that the fusion results transfer more information from the original image and experience less distortion.
+In the qualitative comparison, Figure 5, Figure 6, and Figure 7 present representative samples from the respective datasets.
+Our method demonstrates superior performance in the mean squared error (MSE) graph, indicating that the fusion results are closer to the ground truth.
+Compared to other approaches, our method achieves more accurate restoration of both spectral and spatial details, underscoring the effectiveness of our fusion technique.
+To further verify the generalization ability of our method in full-resolution scenes, we use three non-reference metrics, including ${D}_{s}$}, ${D}_{\lambda}$}, and QNR, to evaluate our method on the full WorldView-II dataset.
+The metric results are summarized in Table II, while Figure 8 provides a visual comparison.
+Our method consistently outperforms the competing approaches across all metrics, highlighting the robust adaptability of Shuffle Mamba in image fusion tasks.
+To comprehensively evaluate the parameter efficiency and computational performance of the proposed method, we further report additional metrics, including the number of parameters, GFLOPs, and actual inference time.
+All methods were benchmarked on $128\times 128$ resolution images from the WorldView-II dataset, and the results are summarized in Table III.
+Our method achieves a comparable parameter count to recent state-of-the-art (SOTA) approaches, while being significantly more lightweight—requiring only one-third to one-half the parameters of FAME [13] and DISPNet [52].
+Moreover, it consistently outperforms the latest ARConv [54] method in terms of parameter count, computational cost, and inference time.
+As for training efficiency, due to substantial differences in model structures, we provide a focused comparison between Pan-mamba [12] and our method under the same number of epochs and settings.
+Pan-mamba completes training in 4.6301 hours (100%), while our method takes 5.2194 hours (113%).
+This indicates a modest 13% increase in training time.
+However, our method achieves consistently superior performance across both quantitative and qualitative evaluations, demonstrating its effectiveness in enhancing image fusion quality.
+Overall, these results highlight the efficiency and practicality of the proposed Shuffle Mamba framework, offering a compelling trade-off between performance and computational cost.
+IV-C2 Medical Image Fusion
+The quantitative comparison results for four metrics on the MIF dataset are presented in Table IV.
+The proposed method demonstrates strong performance across nearly all metrics, highlighting its effectiveness in medical image fusion.
+In our experimental results, a higher VIF score indicates closer alignment with human visual perception.
+Enhanced SCD, Qabf, and SSIM scores suggest that the fused image preserves greater similarity and undergoes less distortion compared to the original images.
+Qualitative comparisons of several methods on the MRI datasets are shown in Figure 9, Figure 10, and Figure 11.
+Our method consistently delivers superior visual quality, a finding further supported by the experimental metrics.
+Furthermore, to evaluate perceptual quality, we performed a user study involving ten participants with medical backgrounds, including graduate students and clinical researchers.
+Each of the 72 MIF test pairs (ours vs.
+CDDFuse) was randomly displayed for pairwise comparison.
+The proposed method was preferred in 60 out of 72 cases (83.3%), primarily due to clearer anatomical boundaries and improved soft-tissue visibility, supporting the quantitative superiority with perceptual evidence.
+IV-C3 Infrared and Visible Image Fusion
+To further evaluate the generalization capability of the proposed method across different multi-modal image fusion tasks, we additionally conducted experiments on Infrared and Visible Image Fusion (IVIF).
+In the IVIF setting, all methods were trained on the MSRS dataset [49] and tested on three benchmark datasets: MSRS, RoadScene [61], and M3FD [31].
+Besides the comparison methods used in the MIF experiments, including U2Fusion [60], SwinFusion [36], CDDFuse [71], PSLPT [53], FusionMamba [43], and TC-MoA [77], we further extended the comparison to include SemLA [59], MLFuse [21], and Up-Fusion [23].
+Table V presents the quantitative results on three datasets.
+It can be observed that the proposed method consistently achieves the best overall performance across all evaluation metrics and datasets.
+Specifically, on the MSRS dataset, our method attains the highest scores in SCD, VIF, Qabf, and SSIM, indicating superior capability in preserving complementary modal information and structural consistency.
+For the RoadScene dataset, our method ranks first on three metrics and achieves the second-best performance on Qabf, demonstrating robust generalization to complex driving scenarios with varying illumination and scene dynamics.
+These results validate that the proposed framework not only excels in the pan-sharpening and MIF but also exhibits strong cross-modal generalization and adaptability in the IVIF task.
+IV-D Ablation Experiments
+To evaluate the effectiveness of the Random Shuffle mechanism, we conducted ablation experiments by removing the shuffle operation from the RM, RCIM, and RMIM modules, individually and in combination.
+As shown in Table VI, removing the shuffle operation consistently leads to performance degradation across all three datasets, confirming that the proposed scanning strategy is a key contributor to fusion quality.
+Notably, the degree of performance drop varies across datasets and modules.
+On WorldView-II, removing the shuffle in the RCIM module results in the largest decline.
+When multiple shuffle operations are removed simultaneously, the performance loss becomes more pronounced, demonstrating that these modules provide complementary gains.
+These observations confirm that the random-shuffle scanning strategy enhances global feature diversity and effectively reduces modality bias in different fusion stages.
+In addition, we further evaluated the influence of different scanning strategies through a comparative ablation study, as presented in Table VII.
+In this experiment, the default Random Shuffle Scanning (RSS) was replaced with Sequential Scanning (SS), Bidirectional Scanning (BS), and Diagonal Scanning (DS), respectively.
+The results show that RSS consistently outperforms all deterministic scanning strategies in all metrics.
+Specifically, while deterministic strategies such as BS and DS yield marginal improvements over SS, none of them can match the performance achieved by RSS.
+This confirms that introducing stochasticity through the Random Shuffle mechanism enables more diverse and global spatial interactions, effectively improves global context aggregation, and ultimately leads to superior fusion performance.
+Figure 12: The effective receptive fields of different scanning schemes.
+To validate the effectiveness of the proposed random shuffle scanning strategy, we provide a comparative visualization of the effective receptive fields (ERFs) under different scanning schemes.
+As illustrated in Figure 12, the long-range dependencies captured by 2D scanning are primarily concentrated along horizontal and vertical directions.
+At the same time, diagonal and zigzag scans exhibit global receptive fields aligned with their respective diagonal paths.
+In contrast, our method produces a more uniformly distributed ERF, characterized by overall darker responses without prominent directional preferences.
+This demonstrates that the proposed random shuffle strategy effectively mitigates the local bias inherent in fixed scanning patterns, facilitating more balanced global modeling.
+Figure 13: The performance and costs in testing time and memory when choosing different numbers of samples.
 Due to Monte Carlo averaging, Shuffle Mamba makes trading memory and computational resources possible for improved performance.
 We studied the relationship between the number of samples, the PSNR index, and resource consumption.
-We experimented five times for each sample size, calculating the average and standard deviation of diverse metrics.
-Figure 7 (a) illustrates the PSNR trend, while Figure 7 (b) and (c) details the corresponding memory usage and the time required to process each image.
+We conducted five experiments for each sample size, calculating the average and standard deviation of various metrics.
+Figure 13 (a) illustrates the PSNR trend, while Figure 13 (b) and (c) detail the corresponding memory usage and the time required to process each image.
 As the number of samples increases, performance and resource consumption rise, enabling a trade-off between performance and efficiency.
 Additionally, Monte Carlo averaging enhances the theoretical robustness of the random shuffle operation, effectively improving the mean of PSNR while reducing the variance of the fusion result.
-More experimental results can be found in the supplementary materials.
-Figure 7: The performance and costs in testing time and memory when choosing different numbers of samples.
-Conclusion
+It is noteworthy that even when the number of Monte Carlo samples is set to 1, the network still performs well.
+This is because the overall architecture consists of multiple stacked modules, each with independent random shuffle scans.
+Therefore, multiple instances of random shuffling and recovery occur within the network, which further ensures stable performance and demonstrates the robustness of the proposed method.
+V Conclusion
 In this paper, we replace the fixed scanning strategy used in current Mamba-based methods with a random shuffle-based scanning method and design a new Shuffle Mamba framework.
-This approach mitigates the bias introduced by fixed scanning methods and provides an global receptive field for multi-modal image fusion.
-During testing, we employ Monte Carlo averaging to account for the introduced random factors.
+This approach mitigates the bias introduced by fixed scanning methods and provides a global receptive field for multi-modal image fusion.
+During testing, we employ Monte Carlo averaging to account for the random factors introduced.
 Extensive experiments on two tasks demonstrate that our approach outperforms state-of-the-art methods and exhibits strong generalization capabilities.
+Despite these advantages, our method has certain limitations.
+The primary drawback lies in the computational cost associated with the Monte-Carlo averaging during inference.
+As analyzed in Figure 13, while increasing the number of scan samples improves the results and reduces variance, it inevitably increases inference time and memory consumption linearly.
+Achieving the theoretical upper bound of performance requires multiple forward passes, which may limit real-time applications on resource-constrained edge devices.
+In future work, we aim to explore more efficient scanning strategies that maintain unbiased global perception without requiring repetitive sampling.
+Additionally, inspired by recent trends, we plan to extend the Shuffle Mamba framework to handle complex fusion scenarios, such as misaligned inputs and extreme weather conditions.